# OpenEarthAgent[Uncaptioned image]: A Unified Framework for Tool-Augmented Geospatial Agents

### v1

@@ 6 Conclusion
Figure A1 presents a zero-shot example demonstrating the geospatial reasoning capability of OpenEarthAgent. Given a natural language instruction, the agent autonomously composes a complete multi-step workflow. The agent first formulates a plan to extract the park boundary, compute NBR layers, evaluate their difference, and visualize the result. It sequentially invokes the tools GetAreaBoundary, AddIndexLayer (for December 2024 and February 2025), ComputeIndexChange, and ShowIndexLayer, generating a differenced NBR (dNBR) layer that quantifies vegetation loss and recovery. The reasoning trace(middle) of Figure A1 illustrates coherent tool grounding and correct argument formatting across the entire chain.
@@
Figure A2 illustrates an example of OpenEarthAgent performing metric-scale geospatial reasoning. Given a natural-language query about the distance between two vehicles with a GSD of 0.072, the agent composes a workflow: detecting vehicles via TextToBbox, computing centroids with Solver, and calculating Euclidean distance using Calculator. The reasoning trace shows coherent tool use and precise cross-modal computation. In Figure A3 the agent extracts the buffered area, adds POI layers, computes nearest-neighbor road distances from each kindergarten to the closest police station, and visualizes the POIs and distance links on a base map. This demonstrates the agent’s ability to integrate spatial query and interpretable visualization within a unified reasoning pipeline.
@@
We conduct a fine-grained analysis of model failures to better understand where tool-augmented reasoning breaks down. Figure A4 shows errors across 4 types, and reports their frequency across some evaluated models.
@@
Tool Name Description Input Output Calculator Evaluates math expressions using Python. Math expression (Python syntax) Numeric result as text OCR Extract text from an image. Image Text with bounding boxes DrawBox Draws a box on an image. Image, box coords, optional label Image with drawn box AddText Adds text to an image. Image, text, position, optional color Image with text overlay GoogleSearch Returns search results from Google. Query string, optional number of results Search result text Plot Plots data using Python code. Python code defining solution() Generated plot image Solver Solves equations using sympy. Python code defining solution() and using sympy Equation solution as string TextToBbox Finds object based on description. Image, object description Bounding box coordinates ImageDescription Generates caption for image. Image Text description RegionAttribute- Description Describes attribute in region or image. Image, attribute, optional bbox Text description of attribute CountGivenObject Counts objects in an image. Image, object name, optional bbox Integer count ChangeDetection Describes changes between two images. Text query, pre and post images Description of changes SegmentObject- Pixels Segments objects and counts pixels. Image, object name, optional flag Pixel count(s) ObjectDetection Detects objects in the image. Image Labels, boxes, scores GetAreaBoundary Gets area boundary as GeoPackage. Place name or bbox, optional buffer GeoPackage with boundary AddPoisLayer Adds POIs to GeoPackage. GeoPackage, query, layer name Confirmation with POI count ComputeDistance Computes distance between layers. GeoPackage, source and target layers Summary text and layer name DisplayOnMap Renders the map image from layers. GeoPackage, layer names PNG image path AddIndexLayer Computes spectral index layer. GeoPackage, index type, year, name Saved layer and class stats ComputeIndex- Change Computes change between index layers. GeoPackage, index type, layer names Saved change layer and stats ShowIndexLayer Generates a preview of the raster index. GeoPackage, index type, layer name Preview image path GetBboxFrom- Geotiff Extract an area bounding box (W, S, E, N) from a GeoTIFF GeoTiff GeoPackage with boundary DisplayOnGeotiff Render GeoPackage layers (with feature names) over a given GeoTIFF GeoPackage, layers, GeoTiff rendered GeoTiff Terminate Ends task and returns final answer. Final answer text None

### v2

@@ 5 Conclusion
Fig. A1 presents a zero-shot example demonstrating the geospatial reasoning capability of OpenEarthAgent. Given a natural language instruction, the agent autonomously composes a complete multi-step workflow. The agent first formulates a plan to extract the park boundary, compute NBR layers, evaluate their difference, and visualize the result. It sequentially invokes the tools GetAreaBoundary, AddIndexLayer (for December 2024 and February 2025), ComputeIndexChange, and ShowIndexLayer, generating a differenced NBR (dNBR) layer that quantifies vegetation loss and recovery. The reasoning trace(middle) of Fig. A1 illustrates coherent tool grounding and correct argument formatting across the entire chain.
@@
Fig. A2 illustrates an example of OpenEarthAgent performing metric-scale geospatial reasoning. Given a natural-language query about the distance between two vehicles with a GSD of 0.072, the agent composes a workflow: detecting vehicles via TextToBbox, computing centroids with Solver, and calculating Euclidean distance using Calculator. The reasoning trace shows coherent tool use and precise cross-modal computation. In Fig. A3 the agent extracts the buffered area, adds POI layers, computes nearest-neighbor road distances from each kindergarten to the closest police station, and visualizes the POIs and distance links on a base map. This demonstrates the agent’s ability to integrate spatial query and interpretable visualization within a unified reasoning pipeline.
@@
We conduct a fine-grained analysis of model failures to better understand where tool-augmented reasoning breaks down. Fig. A4 shows errors across 4 types, and reports their frequency across some evaluated models.
@@
Tool Name Description Input Output Calculator Evaluates math expressions using Python. Math expression (Python syntax) Numeric result as text OCR Extracts text from an image. Image Text with bounding boxes DrawBox Draws a box on an image. Image; box coordinates; optional label Image with drawn box AddText Adds text to an image. Image; text; position; optional color Image with text overlay GoogleSearch Returns search results from Google. Query string; optional number of results Search result text Plot Plots data using Python code. Python code defining solution() Generated plot image Solver Solves equations using sympy. Python code defining solution() and using sympy Equation solution as string TextToBbox Finds object based on description. Image; object description Bounding box coordinates ImageDescription Generates a caption for an image. Image Text description RegionAttribute - Description Describes an attribute in a region or image. Image; attribute; optional bounding box Text description of attribute CountGivenObject Counts objects in an image. Image; object name; optional bounding box Integer count ChangeDetection Describes changes between two images. Text query; pre- and post-images Description of changes SegmentObject - Pixels Segments objects and counts pixels. Image; object name; optional flag Pixel count(s) ObjectDetection Detects objects in the image. Image Labels; boxes; scores GetAreaBoundary Gets area boundary as a GeoPackage. Place name or bounding box; optional buffer GeoPackage with boundary AddPoisLayer Adds POIs to a GeoPackage. GeoPackage; query; layer name Confirmation with POI count ComputeDistance Computes distance between layers. GeoPackage; source layer; target layer Summary text and layer name DisplayOnMap Renders the map image from layers. GeoPackage; layer names PNG image path AddIndexLayer Computes spectral index layer. GeoPackage; index type; year; name Saved layer and class statistics ComputeIndex - Change Computes change between index layers. GeoPackage; index type; layer names Saved change layer and statistics ShowIndexLayer Generates a preview of the raster index. GeoPackage; index type; layer name Preview image path GetBboxFrom - Geotiff Extracts an area bounding box (W,S,E,N) from a GeoTIFF. GeoTIFF GeoPackage with boundary DisplayOnGeotiff Renders GeoPackage layers (with feature names) over a given GeoTIFF. GeoPackage; layers; GeoTIFF Rendered GeoTIFF Terminate Ends task and returns final answer. Final answer text None